A study on the delivery time of mail between two locations has found that the data follows a normal distribution. The study has also found that on average, mail takes up to five days to arrive with a standard deviation of 0:75 days. Using this information, what is the probability that the mail will take between four and six days to arrive?
Answer by Boreal(15235)

z=(x-mean)/sd
z(4)=(4-5)/0.75=(-1/.75)=-1.33
z(6)=1/0.75=+1.33
probability of z between these two is 0.8165